A traveling wave is given by the expression y = 0.30 sin (0.49πx + 0.90πt) m . What is the group velocity of the wave? What is the instantaneous displacement of the wave at x = 0.7 m and t = 0.2 s
Answer #1

wave velocity

v = w/k = 0.9* 3.14 / 0.4* 3.14 = 2.25 m/s

instantaneous position

y = 0.3 sin ( 0.49 * 3.14 * 0.7 + 0.9* 3.14 * 0.2)

y = 0.2992 m

========

Comment before rate in case any doubt, will reply for sure.. goodluck